OpenStudy (anonymous):

Which of the following equations has the solution set {0}? 1)-14m - 7m + 1 = -7m + 1 2)-14m + 7m + 1 = -7m + 2 3)14m - 7m - 1 = -7m + 1

OpenStudy (texaschic101):

-14m - 7m + 1 = -7m + 1 -21m + 1 = -7m + 1 --- add 7m to both sides -21m + 1 = 1 -- subtract 1 from both sides -21m = 0 m = 0/-21 m = 0